OpenStudy (anonymous):

If I want to find the Radius, Height, and Diameter of a cone and it already has the volume given at 60pi, how would I do that?

OpenStudy (anonymous):

@iGreen

OpenStudy (tkhunny):

There are infinitely many such cones. You will need more information.

OpenStudy (jhannybean):

think about the volume of a cone: \(\large V = \pi r^2 \left(\frac{h}{3}\right)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Ah, choices! Well, that does narrow it down, but all those choices work. We'll still need more information. Choices for what? Are those r, d, or h?

OpenStudy (anonymous):

Out of all of those choices 3 of them go into r, h, and d

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@amistre64 @iGreen @EclipsedStar